A HUMAN RESOURCE MANAGER IS CONSIDERING REPLACING TRADITIONAL PROGRAM WITH A NEW VIRTUAL TRAINING PROGRAM FOR NEW 10 EMPLOYEES THE SALARY OF THE HUMAN RESOURCE STAFF IS $20 PER HOUR, AND SALARY OF NEW EMPLOYEES IS $12.50 PER HOUR. ASSUMING THE QUALITY OF PROGRAM IS EQUAL, WHICH THE MOST COST EFFICIENT?
(PROGRAM A: PROGRAM COST $14,000 ;ADMINISTRATION FEE $145 ;HR STAFF NEEDED 3 ; HOURS TO COMPLETE TRAINING 25)
(PROGRAM B: PROGRAM COST $13,000 ; ADMINSTRATION FEE $170 ;HR STAFF NEEDED 4 ;HOURS TO COMPLETE TRAINING 28)
(PROGRAM C : PROGRAM COST $14,500 ;ADMINSTRATION FEE $140 ; HR STAFF NEEDED 4 ; HOURS TO COMPLETE TRAINING 25)
(PROGRAM D : PROGRAM COST $16,000 ; ADMINSTRATION FEE $130 ; HR STAFF NEEDED 3 ; HOURS TO COMPLETE TRAINING 30)
1
Expert's answer
2020-03-23T15:50:58-0400

For the first program:

"14000 + 145 + 3 \times 20\times25 + 10 \times 12.5 \times25 = 18770"

For the second program:

"13000 + 170 + 4 \times 20\times 28 + 10 \times12.5\times 28 = 18910"

The third:

"14500 + 140 + 4\times20\times25 + 10\times12.5\times 25 = 19765"

The fourth:

"16000 + 130 + 3 \times20\times30 + 10\times12.5\times30 = 21680"

 THE MOST COST EFFICIENT is program A